OpenStudy (anonymous):

how to factor a polynomial to find zeros and multiplicity. polynomial is x^4-7x^3+12x^2+4x-16

myininaya (myininaya):

try factors of -16 to see if you get remainder 0 or you could plug into the function and see if the output is 0 if the output is 0 then that number that you plugged in is a zero

OpenStudy (anonymous):

use the factor theorem eg try f(1) - if this is 0 then x-1 is a factor: f(1) = 1 - 7 + 12 + 4 -16 = -6 - no)) f(-1) = 1 + 7 + 12 - 4 - 16 = 0 - so (x+ 1) is a factor now divide the polynomial by x+ 1 to give a cubic expression and comtinue as above

myininaya (myininaya):

lets try -4 -4 | 1 -7 12 4 -16 | -4 -44 ------------------- 1 -11 -32 so this is defintely not gonna give us remainder 0 lets try x=-2 -2| 1 -7 12 4 -16 | -2 18 ----------------- 1 -9 30 this will not work either lets try x=1 1 | 1 -7 12 4 -16 | 1 -6 6 10 -------------------- 1 -6 6 10| -6 this will not work either since remainder is -6 try x=-1 -1 | 1 -7 12 4 -16\ | -1 8 -20 16 -------------------------- 1 -8 20 -16 | 0 so x=-1 works which means x+1 is a factor and x^3-8x^2+20x-16 is also a factor

OpenStudy (anonymous):

to give a trinomial which you may be able to factor

myininaya (myininaya):

lets try x=2 2 | 1 -8 20 -16 | 2 -12 16 ---- ------------------------- 1 -6 8 | 0 so x=2 works which implies x-2 is also a factor and x^2-6x+8 is another factor

myininaya (myininaya):

you should be able to get the rest

myininaya (myininaya):

since we have x^4-7x^3+12x^2+4x-16=(x+1)(x-2)(x^2-6x+8)

myininaya (myininaya):

so are the factors of x^2-6x+8

myininaya (myininaya):

what*